# Neural Algorithmic Reasoning for Hypergraphs with Looped Transformers

### v2

@@
Xiaoyu Li xiaoyu.li2@student.unsw.edu.au. University of New South Wales.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Jiangxuan Long lungchianghsuan@gmail.com. South China University of Technology.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Zhen Zhuang zhuan244@umn.edu. University of Minnesota.
@@ 3.4 Looped Transformers
Let $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{a}}$ be the weight matrices of query and key, $W_{V}\in\mathbb{R}^{d\times d}$ be the weight matrix of value, and $\sigma$ be the hardmax [^1] The hardmax, a.k.a., $\arg\max$, is defined by $[\sigma(\Phi)]_{i}:=\sum_{k\in K}e_{k}/|K|$, where $e_{k}$ is the standard basis vector for any $k\in[K]$ and $K=\{k~{}|~{}\Phi_{ik}=\max\{\Phi_{i}\}\}.$ function. We define the single-head attention $\psi^{(i)}$ as
@@
$f_{\mathrm{attn}}(X,\widetilde{A})~{}$ $:=~{}\sum_{i\in M_{A}}\psi^{(i)}(X,\widetilde{A})+\sum_{i\in M_{A^{\top}}}\psi^{(i)}(X,\widetilde{A}^{\top})$
$+\sum_{i\in M}\psi^{(i)}(X,I_{n+1})+X,$
where $M_{A}$, $M_{A^{\top}}$, $M$ are the index set of the attention incorporated the incident matrix, and the attention incorporated the transpose of the incident matrix, and attention heads for the standard attention which is defined in [97].
@@
Let $\phi$ be the ReLU function, let $W\in\mathbb{R}^{d\times d}$ be the weight of the MLP, where $d$ is the feature dimension of $X$, let $m$ be the number of layers. For $j=[m]$, we define the MLP layer as
$f_{\mathrm{mlp}}(X):=Z^{(m)}W^{(m)}+X,$
where $Z^{(j+1)}:=\phi(Z^{(j)}W^{(j)})$ and $Z^{(1)}:=X$.
@@
$f(X,\widetilde{A}):=f_{\mathrm{mlp}}(f_{\mathrm{attn}}(X,\widetilde{A})).$
@@
Let $\widetilde{A}$ be defined in Definition 3.3, let $m=O(1)$ denote the layer of transformers. Let the transformer layer $f$ be defined in 3.10. We define the multi-layer transformer as
$h_{T}(X,\widetilde{A}):=f_{m}\circ f_{m-1}\circ\cdot\circ f_{1}(X,\widetilde{A}).$
In the construction of this paper, we set $X\in\mathbb{R}^{K\times d}$, where $K$ is defined in Definition 3.3, and $d$ is a constant independent of $K$. The matrix $X$ stores different variables in its columns. A variable can either be an array or a scalar. Scalars are stored in the top row of the corresponding column, leaving the remaining $K-1$ rows as 00. Arrays are stored in the bottom $K-1$ rows of the column, leaving the top row as 00. We use $B_{\mathrm{global}}$ to represent a column where only the top scalar is $1$, while the rest of the entries are 00. Conversely, we use $B_{\mathrm{local}}$ to represent a column where the top scalar is 00, while the remaining entries are $1$. Additionally, we use $P$ to represent the two columns of position embeddings, where the first column contains $\sin(\theta_{i})$, and the second column contains $\cos(\theta_{i})$ for $i\in[K-1]$. Finally, $P_{\mathrm{cur}}$ is used to represent the position embedding of the current vertex or hyperedge.
@@
Initialize the positional encoding with $p_{0}=\begin{pmatrix}0\\ 1\end{pmatrix}\in\mathbb{R}^{2}$. For each node $i\geq 1$, the positional encoding $p_{i}\in\mathbb{R}^{2}$ is defined recursively by
$p_{i}=R_{\widehat{\delta}}^{\top}p_{i-1}.$
The positional encoding for node $i$ is represented as the tuple $(p_{i}^{(1)},p_{i}^{(2)})$, where $p_{i}^{(1)}$ and $p_{i}^{(2)}$ are the first and second components of the vector $p_{i}$, respectively.
@@ 4.1 Degradation
1:$\mathrm{idx}_{\mathrm{hyperedge}}\leftarrow 0$
2:$\mathrm{val}_{\mathrm{hyperedge}}\leftarrow{\bf 0}_{n_{v}}\|{\bf 1}_{K-1-n_{v}}$
3:$\mathrm{iszero}_{\mathrm{hyperedge}}\leftarrow{\bf 0}_{n_{v}}\|{\bf 1}_{K-1-n_{v}}$
4:$\mathrm{candidates}_{\mathrm{hyperedge}}\leftarrow\Omega\cdot{\bf 1}_{K-1}$
5:$\mathrm{visit}_{\mathrm{hyperedge}}\leftarrow{\bf 0}_{n_{e}}\|{\bf 1}_{K-1-n_{e}}$
6:$\mathrm{termination}_{\mathrm{hyperedge}}\leftarrow 0$triangleright Initialization of visiting hyperedges
@@
13:     $\mathrm{idx}_{\mathrm{hyperedge}}\leftarrow\mathrm{idx}_{\mathrm{hyperedge}}+1$triangleright Increment, Lemma C.2
14:     $\mathrm{val}_{\mathrm{hyperedge}}\leftarrow A[:,\mathrm{idx}_{\mathrm{hyperedge}}]$triangleright Read column from incident matrix, Lemma C.7
@@
16:         $\mathrm{iszero}_{\mathrm{hyperedge}}[i]\leftarrow(\mathrm{val}_{\mathrm{hyperedge}}[i]\leq 0)$triangleright Compare, Lemma C.3
@@
20:              $\mathrm{val}_{\mathrm{hyperedge}}[i]\leftarrow\Omega$triangleright Selection, Lemma C.1
@@
25:              $\mathrm{candidates}[i]\leftarrow\mathrm{val}_{\mathrm{hyperedge}}[i]$triangleright Update variables, Lemma C.1
@@
28:     $\mathrm{visit}_{\mathrm{hyperedge}}[\mathrm{idx}_{\mathrm{hyperedge}}]\leftarrow 1$triangleright Write scalar to column, Lemma C.5
29:     $\mathrm{termination}_{\mathrm{hyperedge}}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit}_{\min})$triangleright Trigger termination, Lemma C.6
30:     $\mathrm{termination}_{\min}\leftarrow\mathrm{termination}_{\min}\land\mathrm{termination}_{\mathrm{hyperedge}}$triangleright AND, Lemma C.8
@@ 5 Key Operation Implementation
Here, we present the selection operation. Given a boolean array, referred to as condition $C$, a target array $E$, and two value arrays, $V_{0}$ and $V_{1}$, the selection operation can achieve the following: when the condition $C$ is $1$, the value from $V_{1}$ will be written to the target array $E$, when the condition $C$ is 00, the value from $V_{0}$ will be written to the target array $E$. By this operation, we can combine two arrays into a single array. Furthermore, by setting the target array $E$ to be the same as one of the value arrays, i.e., $V_{0}$ or $V_{1}$, we can realize conditional updates. This means that the value of an array will be updated if and only if a certain condition holds.
If $V_{0}$, $V_{1}$, $E$, and $C$ are scalars, this operation can be directly applied to scalars, as the lower $[K-1]$ values are kept as 00. In the following lemma, we aim to show that the selection operation can be simulated by a single-layer transformer. See details in Lemma C.1.
@@
Here, we present the operation of increment using positional embeddings. Let $C_{1}$ and $C_{2}$ represent the positional embeddings corresponding to $\sin(\cdot)$ and $\cos(\cdot)$, respectively. The primary objective is to employ a rotation matrix to update the angles encoded in $C_{1}$ and $C_{2}$ by increasing the angular offset $\widehat{\delta}$. The updated values are then written to the target locations $D_{1}$ and $D_{2}$. Typically, when $C_{1}=D_{1}$ and $C_{2}=D_{2}$, this operation can be performed in-place. See details in Lemma C.2.
@@
We present the operation of reading a scalar from a column. Let the target column be denoted as $E$, the position embeddings of the source row $C$ as $C_{1}$ and $C_{2}$, and the source column as $D$. The array is stored in column $D$, and our objective is to extract the scalar at the $C$-th position and write it to the top row of column $E$. In an MLP layer represented as $XW$, we can extract a column using its column index. However, to extract a specific row index, we leverage the property of the positional embedding, $p_{i}^{\top}p_{j}<p_{i}^{\top}p_{i}$ for $i\neq j$. This operation enables more flexible operations on individual values within the array. See details in Lemma C.4.
@@
In the greedy algorithm, we terminate the process and return the result after traversing all objects. Here, we maintain an array, marking the traversed objects as 1. Notably, since the value of $K-1$ does not always equal the number of objects (typically $n_{v}$ or $n_{e}$), we fill the remaining positions with 1. See details in Lemma C.6.
@@
This operation can be applied to both arrays and scalars because, even when the input column is always 00, the result of the AND operation will remain 00. This operation is particularly useful when combining two conditions to trigger a specific operation. See details in Lemma C.8.
@@
Different from the regular AND operation described in Lemma C.8, the repeat AND operation combines a scalar and an array. It can be understood as first replicating the scalar into a length-$K-1$ array, followed by performing the regular AND operation. This operation is often used in nested if statements. See details in Lemma C.9.
@@ 7 Conclusion
Roadmap. In Section A, we introduce more related work. In Section B, we introduce some tools from previous work. In Section C, we present the missing lemma in Section 5. In Section D, we present the missing proof in Section 6. In Section E, we present the missing proof in Section 4.
@@ C.1 Selection
Lemma C.1 (Selection, formal version of Lemma C.1).
@@
$(W^{(1)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(V_{0},V_{0}),(V_{1},V_{1}),(C,C),(E,E),(B_{\mathrm{global}},B_{\mathrm{global}}),(B_{\mathrm{local}},B_{\mathrm{local}})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(2)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(E,E),(V_{0},S_{1}),(V_{1},S_{2})\};\\ -\Omega&\mathrm{if}~{}(a,b)\in\{(C,S_{1}),(B_{\mathrm{global}},S_{2}),(B_{\mathrm{local}},S_{2})\};\\ \Omega&\mathrm{if}~{}(a,b)=(C,S_{2});\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(3)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(E,E),(S_{1},S_{1}),(S_{2},S_{1})\};\\ 0&\mathrm{otherwise,}\end{cases}~{}(W^{(4)})_{a,b}=\begin{cases}1&\mathrm{if}~{}(a,b)=(E,S_{1});\\ -1&\mathrm{if}~{}(a,b)=(E,E);\\ 0&\mathrm{otherwise.},\end{cases}$
@@ C.2 Increment
Lemma C.2 (Increment, formal version of Lemma C.2).
@@
Then we can show that a single-layer transformer can simulate an increment operation, which achieves $X[1,D_{1}]\leftarrow\sin(\arcsin{(X[1,C_{1}])}+\widehat{\delta})$, $X[1,D_{2}]\leftarrow\cos(\arccos{(X[1,C_{2}])}+\widehat{\delta})$.
@@
$(W^{(1)})_{a,b}$ =$=\begin{cases}\cos(\widehat{\delta})&\mathrm{if}~{}(a,b)\in\{(C_{1},S_{1}),(C_{2},S_{2})\};\\ -\sin(\widehat{\delta})&\mathrm{if}~{}(a,b)=(C_{1},S_{2});\\ \sin(\widehat{\delta})&\mathrm{if}~{}(a,b)=(C_{2},S_{1});\\ 1&\mathrm{if}~{}(a,b)\in\{(D_{1},D_{1}),(D_{2},D_{2})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(2,3)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(D_{1},D_{1}),(D_{2},D_{2}),(S_{1},S_{1}),(S_{2},S_{2})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ =$=\begin{cases}1&\mathrm{if}~{}(a,b)\in\{(S_{1},D_{1}),(S_{2},D_{2})\};\\ -1&\mathrm{if}~{}(a,b)\in\{(D_{1},D_{1}),(D_{2},D_{2})\};\\ 0&\mathrm{otherwise.}\end{cases}$

### v3

@@
Zekai Huang and Yingyu Liang and Zhenmei Shi and Zhao Song and Zhen Zhuang The Ohio State University.The University of Hong Kong.University of Wisconsin-Madison. magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.University of Minnesota.
@@ 3.4 Looped Transformers
Let $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{a}}$ be the weight matrices of query and key, $W_{V}\in\mathbb{R}^{d\times d}$ be the weight matrix of value, and $\sigma$ be the hardmax [^1] The hardmax, a.k.a., $\arg\max$, is defined by $\sigma(\Phi)]_{i}:=\sum_{k\in K}e_{k}/|K|$, where $e_{k}$ is the standard basis vector for any $k\in[K]$ and $K=\{k~|~\Phi_{ik}=\max\{\Phi_{i}\}\}.$ function. We define the single-head attention $\psi^{(i)}$ as
@@
[…]
where $M_{A}$, $_{A^{\top}}, MM$ are the index set of the attention incorporated the incident matrix, and the attention incorporated the transpose of the incident matrix, and attention heads for the standard attention which is defined in [97].
@@
Let $\phi$ be the ReLU function, let $W\in\mathbb{R}^{d\times d}$ be the weight of the MLP, where $d$ is the feature dimension of $X$, let $m$ be the number of layers. For $j=[m]$, we define the MLP layer as $f_{\mathrm{mlp}}(X):=Z^{(m)}W^{(m)}+X,$ where $Z^{(j+1)}:=\phi(Z^{(j)}W^{(j)})$ and $Z^{(1)}:=X$.
@@
$).\displaystyle f(X,\widetilde{A}):=f_{\mathrm{mlp}}(f_{\mathrm{attn}}(X,\widetilde{A}$)).
@@
[…]
In the construction of this paper, we set $X\in\mathbb{R}^{K\times d}$, where $K$ is defined in Definition 3.3, and $d$ is a constant independent of $K$. The matrix $X$ stores different variables in its columns. A variable can either be an array or a scalar. Scalars are stored in the top row of the corresponding column, leaving the remaining K1K-1 rows as 0. Arrays are stored in the bottom K1K-1 rows of the column, leaving the top row as 0. We use $B_{\mathrm{global}}$ to represent a column where only the top scalar is $1$, while the rest of the entries are 0. Conversely, we use $B_{\mathrm{local}}$ to represent a column where the top scalar is 0, while the remaining entries are $1$. Additionally, we use $P$ to represent the two columns of position embeddings, where the first column contains $\sin(\theta_{i})$, and the second column contains $\cos(\theta_{i})$ for $i\in[K-1]$. Finally, $P_{\mathrm{cur}}$ is used to represent the position embedding of the current vertex or hyperedge.
@@
Initialize the positional encoding with $p_{0}=\begin{pmatrix}0\\ 1\end{pmatrix}\in\mathbb{R}^{2}$. For each node $i\geq 1$, the positional encoding $p_{i}\in\mathbb{R}^{2}$ is defined recursively by $p_{i}=R_{\widehat{\delta}}^{\top}p_{i-1}.$ The positional encoding for node $i$ is represented as the tuple $(p_{i}^{(1)},p_{i}^{(2)})$, where $p_{i}^{(1)}$ and $p_{i}^{(2)}$ are the first and second components of the vector $p_{i}$, respectively.
@@ 4 Main Results
Section 4.1 resents a dynamic degradation mechanism to extend Dijkstra, BFS, and DFS from graphs to hypergraphs. In Section 4.2, we reformulate the Helly property test into a looped-Transformer-executable algorithm and proves its correctness for weighted hypergraphs. Moving on to Section 4.3, we show the looped Transformer can simulate deterministic hypergraph motif algorithms.
@@ 4.1 Degradation
[…]
@@
13:  $\mathrm{idx}_{\mathrm{hyperedge}}\leftarrow\mathrm{idx}_{\mathrm{hyperedge}}+1$ $\triangleright$ Increment, Lemma C.2
14:  $\mathrm{val}_{\mathrm{hyperedge}}\leftarrow A[:,\mathrm{idx}_{\mathrm{hyperedge}}]$ $\triangleright$ Read column from incident matrix, Lemma C.7
@@
16:   $\mathrm{iszero}_{\mathrm{hyperedge}}[i]\leftarrow(\mathrm{val}_{\mathrm{hyperedge}}[i]\leq 0)$ $\triangleright$ Compare, Lemma C.3
@@
20:     $\mathrm{val}_{\mathrm{hyperedge}}[i]\leftarrow\Omega$ $\triangleright$ Selection, Lemma C.1
@@
25:     $\mathrm{candidates}[i]\leftarrow\mathrm{val}_{\mathrm{hyperedge}}[i]$ $\triangleright$ Update variables, Lemma C.1
@@
28:  $\mathrm{visit}_{\mathrm{hyperedge}}[\mathrm{idx}_{\mathrm{hyperedge}}]\leftarrow 1$ $\triangleright$ Write scalar to column, Lemma C.5
29:  $\mathrm{termination}_{\mathrm{hyperedge}}\leftarrow\neg(0~\mathrm{in}~\mathrm{visit}_{\min})$ $\triangleright$ Trigger termination, Lemma C.6
30:  $\mathrm{termination}_{\min}\leftarrow\mathrm{termination}_{\min}\land\mathrm{termination}_{\mathrm{hyperedge}}$ $\triangleright$ AND, Lemma C.8
@@ 5 Key Operation Implementation
Here, we present the selection operation. Given a boolean array, referred to as condition $C$, a target array $E$, and two value arrays, $V_{0}$ and $V_{1}$, the selection operation can achieve the following: when the condition $C$ is $1$, the value from $V_{1}$ will be written to the target array $E$, when the condition $C$ is 0, the value from $V_{0}$ will be written to the target array $E$. By this operation, we can combine two arrays into a single array. Furthermore, by setting the target array $E$ to be the same as one of the value arrays, i.e., $V_{0}$ or $V_{1}$, we can realize conditional updates. This means that the value of an array will be updated if and only if a certain condition holds.
If $V_{0}$, $V_{1}$, $E$, and $C$ are scalars, this operation can be directly applied to scalars, as the lower [K1][K-1] values are kept as 0. In the following lemma, we aim to show that the selection operation can be simulated by a single-layer transformer. See details in Lemma C.1.
@@
Here, we present the operation of increment using positional embeddings. Let $C_{1}$ and $C_{2}$ represent the positional embeddings corresponding to $\sin(\cdot)$ and $\cos(\cdot)$, respectively. The primary objective is to employ a rotation matrix to update the angles encoded in $C_{1}$ and $C_{2}$ by increasing the angular offset $\widehat{\delta}$. The updated values are then written to the target locations $D_{1}$ and $D_{2}$. Typically, when
$C_{1}=D_{1}$ (1)
and
$C_{2}=D_{2}$ (2)
, this operation can be performed in-place. See details in Lemma C.2.
@@
We present the operation of reading a scalar from a column. Let the target column be denoted as $E$, the position embeddings of the source row $C$ as $C_{1}$ and $C_{2}$, and the source column as $D$. The array is stored in column $D$, and our objective is to extract the scalar at the $C$-th position and write it to the top row of column $E$. In an MLP layer represented as $XW$, we can extract a column using its column index. However, to extract a specific row index, we leverage the property of the positional embedding,
$p_{i}^{\top}p_{j}<p_{i}^{\top}p_{i}$ (3)
for $i\neq j$. This operation enables more flexible operations on individual values within the array. See details in Lemma C.4.
@@
In the greedy algorithm, we terminate the process and return the result after traversing all objects. Here, we maintain an array, marking the traversed objects as 1. Notably, since the value of
$K-1$ (4)
does not always equal the number of objects (typically $n_{v}$ or $n_{e}$), we fill the remaining positions with 1. See details in Lemma C.6.
@@
This operation can be applied to both arrays and scalars because, even when the input column is always 0, the result of the AND operation will remain 0. This operation is particularly useful when combining two conditions to trigger a specific operation. See details in Lemma C.8.
@@
Different from the regular AND operation described in Lemma C.8, the repeat AND operation combines a scalar and an array. It can be understood as first replicating the scalar into a length-K1K-1 array, followed by performing the regular AND operation. This operation is often used in nested if statements. See details in Lemma C.9.
@@ 7 Conclusion
Roadmap. In Section A, we introduce more related work. In Section B, we introduce some tools from previous work. In Section C, we present the missing lemma in Section 5. In Section D, we present the missing proof in Section 6. In Section E, we present the missing proof in Section 4. We discuss our impact statement in Section F.
@@ C.1 Selection
Lemma C.1 (Selection).
@@
$(W^{(1)})_{a,b}$ $=\begin{cases}1&\mathrm{if}~(a,b)\in\{(V_{0},V_{0}),(V_{1},V_{1}),(C,C),(E,E),(B_{\mathrm{global}},B_{\mathrm{global}}),(B_{\mathrm{local}},B_{\mathrm{local}})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(2)})_{a,b}$ $=\begin{cases}1&\mathrm{if}~(a,b)\in\{(E,E),(V_{0},S_{1}),(V_{1},S_{2})\};\\ -\Omega&\mathrm{if}~(a,b)\in\{(C,S_{1}),(B_{\mathrm{global}},S_{2}),(B_{\mathrm{local}},S_{2})\};\\ \Omega&\mathrm{if}~(a,b)=(C,S_{2});\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(3)})_{a,b}$ $=\begin{cases}1&\mathrm{if}~(a,b)\in\{(E,E),(S_{1},S_{1}),(S_{2},S_{1})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ $=\begin{cases}1&\mathrm{if}~(a,b)=(E,S_{1});\\ -1&\mathrm{if}~(a,b)=(E,E);\\ 0&\mathrm{otherwise.},\end{cases}$
@@ C.2 Increment
Lemma C.2 (Increment).
@@
[…]
@@
$(W^{(1)})_{a,b}$ $=\begin{cases}\cos(\widehat{\delta})&\mathrm{if}~(a,b)\in\{(C_{1},S_{1}),(C_{2},S_{2})\};\\ -\sin(\widehat{\delta})&\mathrm{if}~(a,b)=(C_{1},S_{2});\\ \sin(\widehat{\delta})&\mathrm{if}~(a,b)=(C_{2},S_{1});\\ 1&\mathrm{if}~(a,b)\in\{(D_{1},D_{1}),(D_{2},D_{2})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(2,3)})_{a,b}$ $=\begin{cases}1&\mathrm{if}~(a,b)\in\{(D_{1},D_{1}),(D_{2},D_{2}),(S_{1},S_{1}),(S_{2},S_{2})\};\\ 0&\mathrm{otherwise,}\end{cases}$
$(W^{(4)})_{a,b}$ $=\begin{cases}1&\mathrm{if}~(a,b)\in\{(S_{1},D_{1}),(S_{2},D_{2})\};\\ -1&\mathrm{if}~(a,b)\in\{(D_{1},D_{1}),(D_{2},D_{2})\};\\ 0&\mathrm{otherwise.}\end{cases}$